OpenStudy (anonymous):

5m+6m-2m=68 step by step plzz

OpenStudy (anonymous):

5m+6m-2m=68 Combine like terms 9m=68 divide 9 on both sides m=68/9 \[\approx\] 7.6

OpenStudy (agreene):

5m+6m-2m = 68 combine like terms on the left (5+6-2 = 9) 9m=68 divide by 9 m=68/9 or about 7.5555

OpenStudy (anonymous):

combining like terms means combining those terms that have the same variables (i.e. the same letters) in this case, 5m, 6m, and -2m have the same letter, m. Therefore, you can combine them like regular numbers. 5m+6m=11m and 11m-2m is 9m. Dividing both sides by 9 to get m alone becomes 68/9. Hope that helps :)

OpenStudy (anonymous):